OpenStudy (anonymous):

plz help if you can -Simplify 10x/9y − 7 /90 = 3/2y

OpenStudy (xapproachesinfinity):

is this what you have ? \(\Large \frac{10x}{9y}-\frac{7}{90}=\frac{3}{2y}\)

OpenStudy (xapproachesinfinity):

yes or no?

OpenStudy (anonymous):

I got the answer already but thank you